OpenStudy (anonymous):

How do I do this? The following system is solved by substitution. Which expression is substituted for z in the second and third equations? x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

x – 4y – 10 x + 4y – 10 –x + 4y – 10 –x – 4y – 10

OpenStudy (lalaly):

z=-10-x-4y so z=-x-4y-10

OpenStudy (anonymous):

can you show me how? sorry i just want to understand :)

OpenStudy (lalaly):

from the first equation solve for z can you?

OpenStudy (lalaly):

from the first equation x+4y+z=-10 now solve for z z=-10-x-4y

OpenStudy (anonymous):

sorry i didnt get notified that you posted!

OpenStudy (anonymous):

oh wow, i thought it was much more complex... thanks!

OpenStudy (lalaly):

its ok :D im glad i can help